OpenStudy (anonymous):

Indicate the equation of the line through (2,-4) and having slope of 3/5

OpenStudy (anonymous):

use the equation \[(y-y _{1})=m(x-x _{1})\] where m=slope (x1,y1)=(2,-4) in ur problem

OpenStudy (anonymous):

y=mx+b which is y= (3/5)(2x)-4